### v1

# Contextual Prompt Learning for Vision-Language Understanding

Abstract
Recent advances in multimodal learning has resulted in powerful vision-language models, whose representations are generalizable across a variety of downstream tasks. Recently, their generalizability has been further extended by incorporating trainable prompts, borrowed from the natural language processing literature. While such prompt learning techniques have shown impressive results, we identify that these prompts are trained based on global image features which limits itself in two aspects: First, by using global features, these prompts could be focusing less on the discriminative foreground image, resulting in poor generalization to various out-of-distribution test cases. Second, existing work weights all prompts equally whereas our intuition is that these prompts are more specific to the type of the image. We address these issues with as part of our proposed Contextual Prompt Learning (CoPL) framework, capable of aligning the prompts to the localized features of the image.
Our key innovations over earlier works include using local image features as part of the prompt learning process, and more crucially, learning to weight these prompts based on local features that are appropriate for the task at hand. This gives us dynamic prompts that are both aligned to local image features as well as aware of local contextual relationships. Our extensive set of experiments on a variety of standard and few-shot datasets show that our method produces substantially improved performance when compared to the current state of the art methods. We also demonstrate both few-shot and out-of-distribution performance to establish the utility of learning dynamic prompts that are aligned to local image features.
1 Introduction
Figure 1: Left hand side picture describes the arechitecure of the CoCoOp and right hand side highlights the architecture of CoPL. In case of CoCoOp, the global image features considered as the coditioned vector for the prompt tokens where as for our CoPL model, we consider localized image features. The colour scheme highlights the weighted prompts according to their semantic alignment with the local image features.
Fully supervised computer vision models for problems like classification are typically trained on datasets like ImageNet [5], OpenImages [32], JFT300M [29, 57] etc., and have also proven themselves to be effective for a variety of downstream tasks via transfer learning [50, 20, 59]. Despite this, it is challenging to adapt these models to other domains due to various reasons including limited data, annotation constraints etc. Additionally, since these models are trained for specific objectives like classification, they tend to capture concepts related to categories seen during training and not to scale to unseen classes during inference.
To address the finetuning issue above, there has been recent effort in tuning the associated prompts (instead of the model weights). Inspired by traditional prompt engineering efforts [49, 3, 53] , there has been some work in tuning discrete prompts from predefined prompt templates [51] that help in capturing rich semantics from user intents and align them to visual contents. However, since building a rich semantic based prompt templates require domain specific and linguistic knowledge, it is tedious and time consuming job to perform. The CoOp [63] algorithm used ideas from soft-prompting in natural language processing [17, 26, 33, 37, 43, 42] to train dynamic learnable prompt vectors with backpropagation and preserve the semantic relationship between sentences and labels [43]. However, the context lerned with CoOp in this fashion fails to generalize to unseen classes, leading to the need for dynamically updating prompts based on the image context, an idea that was proposed in CoCoOp [62]. This model was trained by explicitly conditioning the prompts on image feature vectors as tokens where a separate lightweight neural network (called meta-net in their work) was used to equally weight all the prompt vectors.
In our work, we identify two key issues with the aforementioned architecture. First, the features obtained from meta-net in CoCoOp [62] are global in nature and hence susceptible to issues like clutter and noise in many few-shot and out-of-distribution test cases (we demonstrate this empirically later on). Next, these features are directly added to all the learned prompt vectors, thus resulting in an equal weighting for each prompt. Consequently, this model is unable to learn which of the prompt vectors are more semantically relevant and contextually meaningful during inference, which makes model less generalizable on unseen classes in a zero-shot settings.
To address the aforementioned problems, we propose a new technique called Contextual Prompt Learning (CoPL). Our key ideas include aligning prompts to local image context, realized with local features, and determining which prompts are more sematically relevant conditioned on such local context. Our insight is that by doing so, we are able to learn a more appropriate weighting of the prompts, unlike equal weighting of prior work above, that is semantically reflective of the actual content of the image under consideration. Inspired by the concept of global attention from the NLP literature [44], during training, we propose to align each local feature vector (e.g., computed from a local image patch) to a set of dynamic soft-prompts using a learned context vector that attends to these prompt vectors. An overview of the methdology is shown in Figure 1. This produces a set of attention weights for the prompt vectors that are semantically aligned to local image regions. This results in CoPL learning more generalizable features as we demonstrate with an extensive set of zero- and few-shot classification evaluation results.
We conduct a comprehensive set of experiments on visual classification on $11$ different datasets (see Section 4 and scenarios (zero-shot, one-shot, seen/unseen, and within-dataset and cross-dataset). Across all these experiments, we demonstrate substantial performance improvement when compared to our closest baselines CoOp [63] and CoCoOp [62], indicating the ability of our method to be adapted across various classification settings with little or no training and much reduced prompt engineering.
Our key contributions are summarized below:
  • We identify two key issues with existing prompt-based image classification methods: equal weighting of the prompt vectors and no flow of contextual local information of input images to the prompt vectors while learning during back-propagation.
  • We propose CoPL: Contextualized Prompt Learning, a new method that addresses the issues above by learning prompt weights dynamically and aligning the resulting prompt vectors with local image features.
  • We conduct extensive experiments with CoPL under a variety of classification scenarios and demonstrate substantial performance improvements, in particular in various unseen and few-shot data scenarios. Very interesting to see that, on $11$ different image recognition dataset, on average CoPL achieves state-of-the-art performance on unseen class classifications beating state-of-the-art zeros-hot large-scale model CLIP by $1.4\%$ and conditional prompting model CoCoOp by $3.9\%$ on accuracy. We evaluated CoPL on cross-dataset zero-shot image recognition tasks and on an average on $8$ datasets, CoPL outperformed CoCoOp by $2.3\%$ on accuracy.
2 Related Work
Multimodal Models
Vision-language models have shown great potential in learning generic visual representations. The core idea has been to use natural language supervision for image representation learning and align them jointly in the same embedding space [25, 51]. A vision-language model can be seen as a composition of three parts - a text encoder, an image encoder and a learning methodology that leverages information from both these modalities. Early explorations in this line of research included the problem formulations - metric learning [13], multilabel classification [18, 27], n-gram language learning [34], captioning [8]. Traditionally, hand-crafted descriptors [11, 55] were the mode of capturing image representations. With the introduction deep neural nets, researchers used convolutional neural net based architectures [14, 1] to image representations where as the texts are encoded by simply taking TF-IDF features of the words. With the introduction of pre-trained word vectors the representations learnt by capturing semantic relativity of the words improved [14, 55].
Recent works have focused on learning joint representations of both the modalities using deep learning architectures [15, 25, 38, 28]. As mentioned in [63], there have been significant developments in the areas of transformers[58], and contrastive representation learning [24, 21]. Li et al., [35] proposed VISUALBERT, where texts and images are jointly encoded in a single transformers architecture. The core idea was to align both the modalities inside the transformers architecture by leveraging the cross-attention mapping. The text embddings are obtained from the BERT language model [9] and the image embeddings are obtained from a carefully designed image encoder. One of the biggest release in the multimodal space is the CLIP model [51]. It is a dual encoder based model and during training it matches pairs of images and texts. It is trained on $400$ million training dataset and achieved state-of-the-art performance on different zero-shot downstream tasks. One of the main component of CLIP is the carefully designed prompts which most of the times very hard to formulate. To overcome this, Zhou et al., [63] designed CoOp, which trains dynamic soft-prompts during back-propagation. They keep the underneath multimodal models fixed and only trains the prompt tokens. CoOp has produced impressive results on different image recognition tasks.
Prompting
The intuition behind prompt learning is to capture user intention and instructions to perform certain downstream tasks. In natural language processing literature, capturing instructions to perform various downstream tasks is well established [39, 48, 64]. With the introduction of GPT [3], prompt engineering is shown to be performing efficiently in few-shot knowledge adaptation. Recently Liu et al., [40] designed a template-based question answering model to perform downstream named entity recognition tasks. In their work, they referred to the classical template prompting setup where the templates consist of questions as prompts “What is the [E]?”. In a probabilistic setup, the model predicts the “[E]” token from the distribution of classes which completes the sentence. For example: if the class is “location”, then the model predicts the token and fills in the “[E]” to complete the sentence as “What is the location?”. The key problem behind building prompt templates is to identify the format and linguistic relations with the classes within the dataset. Thus, it requires high skills and knowledge to come up with good prompt templates [41]. Jiang et al., [26] proposed candidate prompts can be generated using text mining, relying on the knowledge of the language model. Deng et al., [7] came up with solution to optimize the discrete prompts using reinforcement learning. The architecture consists of parameter-efficient policy network which generates the discrete prompts after optimizing. They have implemented effective reward stabilization process to tune the rewards efficiently using language models.
Recently, researchers have proposed “soft-prompting” which overcome the shortcomings of the template based prompting including the carefully prompt designing and engineering. The main intuition is to learn dynamic continuous prompt tokens during back-propagation [17, 26, 33, 37, 43, 42]. During optimization, the weights of the language models are fixed. During training, the models predict the classes, carefully mapped from the vocabulary called verbalizers. Recently Goswami et al., [19] in their work highlights the soft prompts can be further tuned with the semantic knowledge of the language models for the different domain adaptation without explicitly verbalizers setup.
Fine-tuning
Before the adaptation to prompt based learning, the standard process was to fine-tune the existing pretrained models on the downstream vision tasks [31, 54, 22]. Gao et al., proposed CLIP-Adapter [16] to improve efficiency of the CLIP by applying a residual transformation layer over the feature embedding or classifier weight. Wortsman et al., [60] proposed WiSE-FT which fine-tunes the layers as a post-ensemble method to improve the efficacy of CLIP.
3 Methodology
Figure 2: The architecture of CoPL: Contextualized Prompt Learning
Here, we introduce and discuss details of our proposed method we call CoPL: Contextualized Prompt Learning. Since CoPL uses the same architectural backbone as CLIP [51], we first begin with a brief review followed by a discussion on how our closest baseline, CoCoOp [62], is trained. This will help set the stage for motivating and discussing our method which we do subsequently. While we use the CLIP backbone for simplicity, we are in no way limited by this design choice. In fact, our method is very much applicable to be used in conjunction with a variety of other architectures, e.g., VisualBERT [35], MDETR [28], GLIP [36] etc.
3.1 Review of CLIP
CLIP [51] is trained using the standard contrastive learning setup where there is a text encoder and an image encoder and the overall objective function is to get their outputs as close as possible in the joint space. The text encoder is implemented with the Transformer model [58] which takes word sequences as input and produces both individual sequence-level as well as overall sentence-level representations. The image encoder is implemented with the ViT architecture [10] which produces local (at the patch level) as well as global image features. The contrastive loss function is designed to capture the similarity between the relevant text and images, that is, the cosine similarity between the related text and image will be maximized whereas the cosine similarity with all other unrelated pairs will be minimized. To be precise, for a $K$-class classification problem, the discrete prompts are designed to have one “class” token. The training objective is then to fill the token with the $i$-th class name where $w_{i}$ is the weight vector generated by the text-encoder for the same.
$p(y|x)=\frac{exp(sim(x,w_{y}))/\gamma}{\sum_{i=1}^{k}exp(sim(x,w_{i}))/\gamma}$ (1)
where $x$ is the image feature vector produced by the image encoder and $\gamma$ is a temperature parameter.
3.2 Review of CoCoOp
CoCoOp: Conditional Context Optimization is built on top of a previously published Context Optimization (CoOp) [63] algorithm. It is observed that the choice of prompts plays a major role in vision-language understanding [51]. However, finding the right match between prompts and description of the image often takes a significant amount of time [63] and sometimes can change the model results drastically. For instance, adding task-specific manual prompts in the Flowers102 dataset [46] would include the word “flowers”, whereas for the DTD dataset [4] the word should be “texture”. In both cases, adding these words in prompts requires manual intervention and careful observation. To overcome such a need for prompt engineering, the CoOp model [63] trains a set of continuous vectors $\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{M}\}$ as context tokens during back-propagation. These vectors are represented as word vectors instead of discrete predefined prompts. Specifically, the prompts for the $i$-th class, denoted by $t_{i}$ can be represented as $[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{M},\mathbf{c}_{i}]$, where $M$ is the number of learnable prompt vectors and $\mathbf{c}_{i}$ is the class embeddings. The CoCoOp [62] algorithm improves CoOp’s [63] performance learning to generate prompts conditioned on each image instance, i.e., these now change for each image unlike in CoOp where they are fixed. CoCoOp does this by training an additional two-layer neural network called meta-net that takes an image feature vector as input and produces a conditional vector that is combined with the prompt vectors to generate the final image-dependent prompts. This is done as follows:
$\mathbf{v}_{m}(\mathbf{x})=\mathbf{v}_{m}+\mathbf{\pi}$ (2)
$\mathbf{\pi}=h_{\theta}(\mathbf{x})$ (3)
where $h_{\theta}$ refers to the meta-net and $\mathbf{x}$ the image feature vector.
3.3 CoPL: Contextual Prompt Learning
While CoCoOp generalizes better than CoOp for unseen-class classification, many issues remain. First, since CoCoOp uses global feature vectors for learning the updated prompts, it focuses less on the discriminative regions in images (which tend to be more local). Next, the addition operation performed in Equation 2 does not capture the individual importance of each prompt token. This is particularly important since certain discriminative regions in images may weight certain prompts more when compared to others and this is not captured in the CoCoOp model. To address these issues, we propose CoPL, a simple and intuitive algorithm that operates at the local feature granularity while also aligning them with prompts to learn prompt importance weights.
Given an image $\mathbf{I}$, we first compute a set of local feature vectors. For instance, this can be the output of a vision transformer model that generates patch embeddings. Let $\mathbf{s}\in\mathbb{R}^{P\times B\times d}$ (where $P$ is number of patches from image, $B$ is training batch size, and $d$ is feature dimesionality) be this. Conditioned on these local features, we determine the semantically most meaningful prompts. To do this, we take inspiration from the attention work of Luong et al. [44] and generate context representations that explicitly consider both the learnable prompt tokens as well as the patch representations. We first learn a lightweight neural network to generate a conditional token for the representation of each patch in $\mathbf{s}$ as:
$\mathbf{s}_{p}=h_{\theta}(\mathbf{s}_{p})$ (4)
where $p\in\{1,2,\ldots,P\}$. This makes the architecture parameter-efficient and easily differentiable during back-propagation. To generate the context representations that can be used to update the prompt tokens, we first learn a variable length alignment vector $\mathbf{a}_{p}$ ($\mathbf{a}\in\mathbb{R}^{B\times M\times d}$), one for every patch $p$ that attends to each prompt tokens $\mathbf{v}_{i}$ and compares them to the corresponding image patch representation $\mathbf{s}_{p}$ as:
$\mathbf{a}_{p}$ $=\text{align}(\mathbf{s}_{p},\mathbf{v}_{i})$ (5)
$=\frac{\text{exp}(\text{score}(\mathbf{s}_{p},\mathbf{v}_{i}))}{\sum_{i=1}^{M}(\text{score}(\mathbf{s}_{p},\mathbf{v}_{i}))}$
where $i\in{1,2,\ldots,M}$ and $M$ is the number of prompt tokens. In our setup, score refers to the content function and is implemented as:
$\text{score}(\mathbf{s}_{p},\mathbf{v}_{i})=\text{tanh}(\mathbf{W}_{a}[\mathbf{s}_{p};\mathbf{v}_{i}])$ (6)
where $W_{a}$ is the weight vector.
Finally, the per-patch context representation $\mathbf{c}_{p}$ is calculated as the weighted sum over all prompt tokens as:
$\mathbf{c}_{p}=\sum_{i=1}^{M}\mathbf{a}_{pi}\mathbf{v}_{i}$ (7)
The final prompt tokens are now obtained by conditioning on the context vectors above as:
$\mathbf{v}_{m}(\mathbf{x})=\sum_{i=1}^{P}\mathbf{v}_{m}+\mathbf{c}_{i}$ (8)
In a nutshell CoPL calculates the prompts for the $i$-th class as $t_{i}=[v_{1}(x),v_{2}(x),...,v_{M}(x),cl_{i}]$, where $cl_{i}$ is the $i$-th class. The prediction probability is calculated as
$p(y|x)=\frac{exp(sim(x,g(t_{y}(x))/\gamma}{\sum_{i=1}^{k}exp(sim(x,g(t_{i}(x))/\gamma}$ (9)
where $g()$ is the feature vector produced by the text encoder. During our entire pipeline, the pre-trained CLIP model is fixed.
4 Experimental Evaluation
Datasets ($\downarrow$) Protocols ($\downarrow$) CLIP CoOp CoCoOp CoPL
Caltech101 Seen 96.8 98.0 97.9 98.1
Unseen 94.0 89.8 93.8 94.9
HM 95.4 93.7 95.8 96.5
OxfordPets Seen 91.1 93.6 95.2 95.6
Unseen 97.2 95.2 97.6 97.8
HM 94.1 94.4 96.4 96.7
StandfordCars Seen 62.3 78.1 70.4 70.7
Unseen 74.8 60.4 73.5 74.4
HM 68.6 68.1 72.0 72.6
Flowers102 Seen 72.0 97.6 94.8 96.1
Unseen 77.8 59.6 71.7 72.1
HM 74.8 74.0 81.7 84.1
Food101 Seen 90.1 88.3 90.7 90.9
Unseen 91.2 82.2 91.2 91.4
HM 90.6 85.1 90.9 91.1
FGVCAircraft Seen 27.1 40.4 33.4 36.1
Unseen 36.2 22.3 23.7 31.3
HM 31.0 28.7 27.7 33.7
SUN397 Seen 69.3 80.6 79.7 80.2
Unseen 75.3 65.8 76.8 77.3
HM 72.2 72.5 78.2 78.7
DTD Seen 53.2 79.4 77.0 78.0
Unseen 59.9 41.1 56.0 50.0
HM 56.3 54.2 64.8 64.0
ImageNet Seen 72.4 76.4 75.9 77.8
Unseen 68.1 67.8 70.4 71.3
HM 70.2 71.9 73.1 74.5
EuroSAT Seen 56.4 92.1 87.4 89.2
Unseen 64.0 54.7 60.0 34.2
HM 60.0 68.9 71.2 61.7
UFC101 Seen 70.5 84.6 82.3 83.1
Unseen 77.5 56.0 73.4 76.6
HM 73.8 67.4 77.6 79.8
Average Seen 69.3 82.7 80.5 81.4
Unseen 74.22 63.2 71.7 75.6
HM 71.7 71.6 75.8 78.5
Table 1: Performance of CLIP, CoOP, CoCoOp, and CoPL on $11$ classification datasets; each of the training dataset consists $16$-shots per class. The dataset highlights CoPL do better generalization than other methods. The baseline results are taken from CoCoOp paper [62]. HM refers to Harmonic Mean between the seen and unseen classes.
Our key insight in CoPL is that we are able to learn generalizable representations by focusing more on the local representations of the image, along with deferentially weighting the learned prompt embeddings. To bring out the efficacy of this, we evaluate our method on multiple settings: $(i)$ knowledge transfer to unseen classes from seen classes within the same dataset, $(ii)$ zero-shot performance across multiple datasets, $(iii)$ efficiency on unseen classes while trained with $1$-shot training set within the same dataset.
Datasets
We follow Zhou et al. [63] to evaluate our model on $11$ image classification dataset of varying complexity. The datasets include: generic classification datasets like ImageNet [6] and Caltech-101 [12]; curated fine-grained datasets like OxfordPets [47], StanfordCars [30], Flowers102 [46], Food101 [2] and FGVCAircraft [45]; scene, action, texture and satellite image recognition datasets from SUN397 [61], UCF101 [56], DTD [4] and EuroSat [23] respectively.
For the few-shot experiments, we follow Zhou et al. [62] to randomly sample datapoints for training. The models are evaluated on the entire test set to report the accuracy. For setting $(i)$ and $(ii)$ (inter-dataset and intra-dataset experiments), we use $16$-shots, consistent with the baseline methods [62]. We use the standard classification accuracy as the evalution metric across all experiments.
Training Details
We build our method on top of the publicly available implementation of CoCoOp [63]. For the CLIP architecture, we use ViT-B/$16$ as image encoder. For all our experiments, we use prompt token length of $4$. All our models are trained with a batch size $1$ for $10$ epochs on a single GPU system. Our models are trained on a $15$ GB Tesla T4444 GPU. We used the standard stochastic gradient descent (SGD) [52] as our optimizer. Our starting learning rate is $0.002$ and we use cosine learning rate scheduler. Our warm-up with a constant learning rate is $0.00001$.
Baseline Models
We compare our approach with the current state-of-the-art method CoCoOp [62], CoOp [63]and also large-scale zero-shot methodology CLIP [51]. While comparing with CLIP, we indeed compare our learned prompt embeddings with manually designed prompts. We closely follow most of our experimental setting with that of Zhou et al., [62].
Methodology
Oxford
Pets
Standford
Cars
Food101 DTD EuroSat Flowers102
FGVC
Aircraft
UCF101 SUN397 Avg
CoCoOp 89.6 61.4 88.2 47.0 63.5 66.6 20.9 65.5 70.3 63.7
CoPL 89.8 61.9 89.5 49.5 58.1 71.3 23.8 70.3 71.9 65.1
Table 2: Zero-shot image recognition results of CoPL across dataset. We trained models on caltech101 dataset and tested on other datasets.
4.1 Knowledge Transfer to Unseen Classes
One of the main issues of CoOp is that it is unable to generalize to the unseen classes while being trained on the base classes. In CoCoOp, the authors claim to bridge the gap by learning prompt conditioned on the global features of the image. But they miss out on capturing the local features of the image and do not give weightage to the semantically aligned learnable prompts, which is the key focus area of our work. Following the implementation of Zhou et al., [62], we conducted our experiments on the above mentioned $11$ datasets both for seen and unseen classes. While training is only conducted on the base classes, during testing we transfer the learnt knowledge to classify unseen classes as well as seen classes.
From Table 1, we can see that across datasets, there is a decrease in performance for CoOp methodology over the zero-shot large-scale method CLIP. Though there is an improvement for the seen classes, the failure across unseen classes makes the model less effective compared to CLIP. Though introduction of conditional prompting improved the performance of CoCoOp, in most of the cases it fails to generalize to unseen classes with the dynamically learnt prompts (in $7$ out of $11$ datasets CLIP has outperformed CoCoOp methodology). This raises the question of how to effectively learn the prompts which are more generalizable and align with the input image features?
Interestingly, our proposed approach CoPL, improves the accuracy of unseen classes over CoCoOp and CoOp for $10$ different tasks. Moreover, the harmonic mean for all these $10$ tasks are greater than the current state-of-the-art model, which specifies that the learned prompts are more generalizable across domains and tasks. It is important to note that CoPL comfortably outperforms manual prompt based methodology CLIP on $5$ different tasks. With the relatively complex dataset like EuroSAT, where localization is hard to be aligned due to the nature of satellite imagery, for the seen classes CoPL outperformed CoCoOp by $1.8\%$ and CLIP by $32.8\%$ on accuracy.
Figure 3: Performance comparison of CoPL with CoOp and CoCoOp for the unseen class detection while trained on base classes within the dataset.
In Figure 3, we further analyse the absolute performance gain on the unseen classes over CoOp and CoCoOp methodologies. Interestingly, when compared with CoOp, for the dataset UFC101, CoPL improve the accuracy by $20.6\%$, enhancing the capability of the dynamic prompt vectors to learn semantic relatedness conditioned on the local image features. For the FGVCAircraft dataset, the absolute performance gain of $9.0\%$ and $7.6\%$ for CoOp and CoCoOp respectively highlights the efficacy of the model. Performance boost of $8.9\%$ on the DTD dataset, confirms the capability of COPL in identifying different textures, opening up the possibility to deploy the model for design understanding. The average performance gain of $2.7\%$ on the FGVCAircraft dataset over CLIP, defines the ability of CoPL to transfer the learnt knowledge to the relatively complex unseen classes.
In general, while considering both seen and unseen dataset, CoPL outperforms CoCoOp by $2.7\%$ in accuracy. At the same time, it achieves a gain of $6.8\%$ on accuracy over manual prompt base methodology CLIP, suggesting that conditional alignment of prompts to local features of the image can generalize the model better to diverse recognition tasks. Ideally, the manually designed prompts should have been well aligned with the images, thus CLIP can be considered as the best performing model. It is interesting to note that CoPL, on an average across all $11$ datasets, achieves state-of-the-art performance by outperforming CLIP by $1.4\%$ in accuracy. This justifies our argument of making dynamic prompt vectors more contextualized like human annotated prompts by observing the different local features of the image.
4.2 Inter Dataset Zero-shot Performance
Here, we test the mettle of our approach on an even more challenging setting. We learn the model on a dataset, and evaluate the model on a different dataset to see how transferable the learned representations are.
Concretely, we train the model on Caltech101 dataset and evalute its performance on the rest of the datasets. Observe, in Table 2, CoPL outperforms CoCoOp across most of the datasets. For complex dataset like FGVCAircraft, CoPL gains $2.9\%$ more accuracy without being explicitly trained on this dataset. As the Caltech101 is a general purpose object classification dataset, we observe impressive performance on OxfordPets and Food101 dataset. Interestingly, though trained on object classification dataset, CoPL is able to transfer the knowledge to do texture recognition task on DTD dataset, outperforming CoCoOp by $2.5\%$ on accuracy. In the most challenging EuroSAT dataset, which has very different image sets (consists of images taken from satellite), CoPL achieves $58.1\%$ accuracy which is comparable to the performance of the CoCoOp. These highlights that CoPL able to learn the semantics of the images by identifying the localized features and align it with the learnable textual prompts.
4.3 One-shot Training
Next, we evaluate our approach in an extremely low data regime. Here, we train CoPL and CoCoOp with $1$ training instance per class for each of the image recognition task and test the accuracy on both seen and unseen classes within the dataset. In our experiments, we valuate across $7$ datasets, and discuss their performance on the seen and unseen classes next.
Figure 4: Performance comparison of CoPL with CoCoOp for the seen class classification while trained on $1$-shot training instance per class.
Seen Class
In Figure 4, we observe that for most of the datasets, CoPL outperforms CoCoOp indicating the capacity of adapting to the in-training datasets. It is interesting to see that, for FGVCAircraft dataset, CoPL outperforms CoCoOp by large margin ($25.9\%$ on accuracy).
Figure 5: Performance comparison of CoPL with CoCoOp for the unseen classes classification while trained on $1$-shot training instance per class.
Unseen Class
The performance of the CoPL on unseen classes while only trained on $1$-shot seen classes is showcased in Figure 5. CoPL outperforms CoCoOp on $6$ datasets. It is interesting to observe that, on FGVCAircraft dataset CoPL improves the accuracy from $4.9\%$ to $28.5\%$.
These results indicates that, the alignment of local image features to prompts helps to capture the semantic meaning of the images while only trained with one training instance per class. This makes the model easily adaptable to diverse set of image recognition task, even for low-resource scenarios.
5 Discussions and Analysis
Local vs Global Image Features
A key contribution of our research work is to utilize the localized image features for prompt learning. Further, we give weightage to the prompts based on the semantic relatedness. In this section, we critically analyse the contribution of using local features an opposed to global features. To evaluate this, we have conduct experiments on Caltech101 and DTD datasets. We selected these two dataset as they are primarily targeted for two different recognition tasks: object recognition and texture recognition. Similar to our earlier evaluation protocol, we evaluate the performance on the seen classes and unseen classes separately.
Methodology Class Type Caltech101 DTD
Global Attention + Global Features Seen 97.2 77.4
Unseen 94.1 39.9
Global Attention + Local Features Seen 98.1 78.0
Unseen 94.9 50.0
Table 3: Ablation Study: Importance of aligning the local contextual image features with the prompt tokens
From Table 3, we can can clearly understand that for both seen and unseen classes, aligning local contextual features with prompts, helps CoPL to generalize better.
Incremental Test
Following the experimental analysis presented in CoCoOp [62], we evaluate the model efficacy where test dataset consists of both seen and unseen classes. In this case, during training the model weights are updated based on the seen classes but has not considered the unseen classes. Thus, during testing model will be performing zero-shot classification on unseen classes.
Methodology Accuracy
CLIP 65.2
CoOp 65.6
CoCoOp 69.1
CoPL 74.7
Table 4: Ablation study: average accuracy on $11$ datasets. The models are trained on training instances for the seen classes. Test datasets are the combination of both seen and unseen classes.
We observe in Table 4, that CoPL comfortably outperforms all the baseline models, suggesting that local image feature alignment with prompts help to generalize across all the classes within a dataset.
Run-time Analysis
We experimentally analyse and quantify the extra wall-clock time that is required for CoPL when compared to the closest baseline CoCoOp. We use Caltech101 for this experiment. For training $800$ data-points (corresponding to $50$ classes) , CoPL took $28$ minutes whereas evaluation on $1549$ data-points took $2$ minutes $05$ seconds. On the other hand CoCoOp takes $21$ minutes $35$ seconds for training and $1$ minute $34$ second for inference. From this, we clearly understand that CoPL significantly improves generalization without sacrificing too much on the computational overhead.
6 Limitation
Our exhaustive experimental analysis across $11$ datasets greatly helped us to test the mettle of our method. Our performance on the EuroSat dataset in the unseen class category, is lower than CoCoOp. While analysing the reasons for the drop in performance, we could uncover that images from EuroSat does not contain images with salient objects. We visualise some such examples in Fig. 6; it can be seen that there are no “local” regions in these images, which can contribute towards modelling the prompt better. This results in lower performance of CoPL on such datasets.
Figure 6: Visualization of samples from EuroSat dataset. We note that majority of the sampled does not contain any salient objects, which makes local feature less effective.
7 Conclusion
In this paper, we present CoPL: Contextual Prompt Learning which can align prompts to the corresponding contextual local image features. During alignment, we also produces a set of attention weights for the prompt vectors that are semantically related to local image regions. Extensive experimental evaluation on $11$ image recognition datasets showcases the efficacy of the methodology in understanding the semantic relationship between the images and the prompts. Moreover, the state-of-the-art zero-shot and few-shot results justify our claim of making CoPL better in generalization by aligning the local features of the images to prompts. The capability of outperforming CLIP in different settings justifies our argument of making dynamic prompt vectors comparable to human-annotated prompts, which opens up the floor to adapt the methodology for the various downstream tasks. In future, we plan to make COPL capable of understanding user intents to make local edits on the images.